i am working on hyperbola's and im not sure if i am understanding it yet. could see if i got the right answer to this one? thanks for any help. this is from a problem we had in class.
The center of the hyperbola 9x^2-16y^2-18x+32y-151=0
the answer i got was (1,1)

Answer by Nate(3500)

9x^2 - 16y^2 - 18x + 32y - 151 = 0
9x^2 - 18x - 16y^2 + 32y = 151
(9x^2 - 18x) + (-16y^2 + 32y) = 151
9(x^2 - 2x) - 16(y^2 - 2y) = 151
9(x - 1)^2 - 16(y - 1)^2 = 151 + 9 - 16 = 144
(x - 1)^2/16 - (y - 1)^2/9 = 1
C(1,1)